OpenStudy (anonymous):

9.55 x 10^24 molecules of CH3OH is how many grams?

OpenStudy (anonymous):

use avogadro's number: \(\sf \color{red}{6.022 \times 10^{22}\frac{mole}{molecule}}\)

OpenStudy (anonymous):

sorry, 23rd power, not 22.

OpenStudy (aaronq):

^he's right. \(\sf moles=\dfrac{molecules}{Avogadro's~number}\)

OpenStudy (anonymous):

Thank you!

OpenStudy (whpalmer4):

It's not \(6.022*10^{23} \dfrac{\text{mole}}{\text{molecule}}\)! It is \(6.022*10^{23}\dfrac{\text{<constituent particle>}}{\text{mole}}\); in this case the constituent particle is a molecule. \[1 \text{ mole} * 6.022*10^{23}\dfrac{\text{molecule}}{\text{mole}} = 1 \cancel{\text{ mole}} * 6.022*10^{23}\dfrac{\text{molecule}}{\cancel{\text{mole}}}\]\[=6.022*10^{23}\text{ molecule}\] Make a point of using the units correctly in all of your work and they will help you find or prevent many errors, especially in problems containing unit conversions.